OpenStudy (anonymous):

solve the proportion. 4/5=x/27 A.21.6 B.20 C.33.75 D.18

OpenStudy (kitten_is_back):

4/5=x/27 5x=108 Cross multiply. Multiply 5 and x. Multiply 4 and 27. x=108/5 Isolate the variable,x by dividing 5 on both sides.